A cheering squad is made up of three members: Lynne of 50kg moving at 3 m/sNorth. Angie of mass 58kg moving eastward at 3 m/s and Mercy of mass 6 kg moving at 3 m/s south. What is the combined momentum of the squad?


1
Expert's answer
2021-12-22T14:09:05-0500

Given:

"m_1=50\:\rm kg"

"m_2=58\:\rm kg"

"m_3=6\:\rm kg"

"v_{1x}=0\:{\rm m/s},\quad v_{1y}=3\:\rm m/s"

"v_{2x}=3\:{\rm m/s},\quad v_{2y}=0\:\rm m/s"

"v_{3x}=0\:{\rm m/s},\quad v_{3y}=-3\:\rm m/s"


The momentum of combined system

"p_x=m_1v_{1x}+m_2v_{2x}+m_3v_{3x}\\
=58*3=174\:\rm kg*m/s"

"p_y=m_1v_{1y}+m_2v_{2y}+m_3v_{3y}\\
=50*3+6*(-3)=132\:\rm kg*m/s"

Magnitude of momentum

"p=\sqrt{p_x^2+p_y^2}=\sqrt{174^2+132^2}=218\:\rm kg*m/s"

Direction of momentum

"\theta=\tan^{-1}\frac{p_y}{p_x}=\tan^{-1}\frac{132}{174}=37^\circ\rm\quad N\:of\: E"